ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 115, 2000

Arias, A. and G. Popescu Noncommutative interpolation and Poisson transforms

205

Bonatti, C. and M. Viana SRB measures for partially hyperbolic systems whose central direction is mostly contracting

157

Chalendar, I., K. Kellay and T. Ransford Binomial sums, moments and invariant subspaces

303

Chen, H. and X. Hua Normal families concerning shared values  355

Globevnik, J. Holomorphic functions which are highly nonintegrable at the boundary

195

Holopainen, I. A sharp L^{q}-Liouville theorem for p-harmonic functions

363

Hua, X. See H. Chen

Iliev, I. D. On the limit cycles available from polynomial perturbations of the Bogdanov–Takens Hamiltonian

269

Ives, D. J. and D. Preiss Not too well differentiable Lipschitz isomorphisms

343

Kellay, K. See I. Chalendar

Kurata, K. On doubling properties for non-negative weak solutions of elliptic and parabolic PDE

285

Lindenstrauss, E. and B. Weiss Mean topological dimension  1

Lomonosov, V. A counterexample to the Bishop–Phelps Theorem in complex spaces

25

Melle-Hernández, A. Milnor numbers for surface singularities  29

Mishchenko, S. P., A. Regev and M. V. Zaicev The exponential growth of codimensions for Capelli identities

333

Popescu, G. See A. Arias

Preiss, D. See D. J. Ives

Ransford, T. See I. Chalendar

Regev, A. See S. P. Mishchenko

Rossmanith, R. Lie centre-by-metabelian group algebras in even characteristic, I

51

Rossmanith, R. Lie centre-by-metabelian group algebras in even characteristic, II 77

Schauenburg, P. Bi-Galois objects over the Taft algebras  101

Sukochev, F. A. Linear topological classification of separable L_{p}-spaces associated with von Neumann algebras of type I

137

Viana, M. See C. Bonatti

Voigt, J. A concentration of mass property for isotropic convex bodies in high dimensions

235

Weiss, B. See E. Lindenstrauss

Yan, J. Oscillations of higher order neutral differential equations of mixed type

125

Zaicev, M. V. See S. P. Mishchenko

Zampieri, G. Solvability of the ¯∂ problem with C^{∞} regularity up to the boundary on wedges of ℂ^{N}

321

Zippin, M. Orthogonal almost locally minimal projections on ℓ_{1}^{n}  253